# Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS–FTIR

### v1

@@ Abstract
This paper presents a photothermal infrared (IR) spectroscopy technique based on a nanoelectromechanical system, which is coupled to a commercial Fourier transform infrared spectrometer (NEMS–FTIR) as a promising solution for the chemical characterization and quantification of nanoplastics. Polystyrene (PS), polypropylene (PP), and polyvinyl chloride (PVC) nanoparticles with nominal diameters of 100, 54, and 262 nm, respectively, were analyzed by NEMS–FTIR with limits of detection (LoD) of 353 pg for PS, 102 pg for PP, and 355 pg for PVC. The PS mass deposited on the NEMS chips was estimated from the measured absorptance values and the attenuation coefficient of PS. The wide spectral range of the FTIR allowed the identification of individual polymer nanoparticles from a mixture. The potential of NEMS–FTIR for the analysis of real–world samples was evaluated by confirming the presence of polyamide (PA) particles released from commercial tea bags during brewing. Accelerated aging of the tea bags under elevated temperature and UV radiation showed continuous release of PA particles over time.
@@ INTRODUCTION
Nanoplastics have become ubiquitous, posing significant challenges linked to their high reactivity, potential to transport pollutants, and widespread dispersion across ecosystems [1, 2, 3, 4, 5, 6, 7]. Their ability to penetrate deep into tissues underscores the need for precise chemical characterization and monitoring [8].
Sampling of nanoplastic particles is particularly challenging due to their nanoscale dimensions and low concentrations, which often exceed the capabilities of many analytical techniques to effectively characterize these tiny fragments [9, 10, 11]. While well–established techniques exist nowadays for microplastic detection and identification, identifying nanoplastics is more challenging.
Surface–enhanced Raman spectroscopy (SERS) has emerged as a promising technique. Zhou et al. [12] and Hu et al. [13] successfully achieved LoDs as low as 5 $\mu$g/mL for PS particles as small as 50 nm. Optical–photothermal IR (O–PTIR) spectroscopy has also proven very effective for the imaging of individual PS nanoparticles as small as 250 nm in mammalian tissues [14] and identification of plastics (> 600 nm) containing polysiloxanes and imides, originating from the steam disinfection of silicone baby teats [15].
@@
Mass spectrometry–based techniques can be used to chemically identify particles smaller than 1 $\mu$m and are often synergistically combined with other methods to reliably detect, quantify, and identify nanoplastics [10, 11, 16]. Pyrolysis–gas chromatography–mass spectrometry (Py–GC/MS) has been extensively applied to detect various types of nanoplastics, including PS, PP, polycarbonate (PC), polyethylene terephthalate (PET), polymethyl methacrylate (PMMA), PVC, polyethylene (PE), and Nylon 6/66, with reported LoDs as low as 1 ng/L [17, 18]. Using the same technique, Li et al. [19] estimated a LoD of 10 ng for two polymer types of nanoplastics in water, PS, and PMMA. A combination of atomic force microscope–IR (AFM–IR) spectroscopy and Py–GC/MS was also used to detect and quantify PE and PVC particles (20–1000 nm), with LoDs of 0.28 $\mu$g for PE and 1.54 $\mu$g for PVC [20]. Alternative approaches, such as thermal desorption–proton transfer reaction–mass spectrometry (TD–PTR–MS), demonstrated remarkable sensitivity, with Materić et al. [21] detecting pure 1 $\mu$m PS particles with an LoD of 0.34 ng, allowing detection of these particles in environmental samples as small as 1 mL without preconcentration.
Many studies have analyzed the release of micro– and nanoplastic particles from tea bags during tea preparation using different analytical methods such as scanning transmission X–ray microscopy (STXM) in combination with near–edge X–ray absorption fine–structure spectroscopy (NEXAFS) [22], IR photothermal heterodyne imaging (IR–PHI) [23], Raman spectroscopy [24], and attenuated total reflectance – FTIR (ATR–FTIR) [25, 26]. These, however, often require sample preconcentration steps or sample preparation procedures which differ from a standard tea brewing procedure.
Although existing analytical techniques offer high sensitivity and specificity, they remain impractical for routine environmental monitoring, particularly for water suppliers and official control laboratories [27]. Mass spectrometry–based techniques provide excellent precision, but their complexity and cost hinder their use in routine environmental monitoring [11, 16, 28]. SERS suffers from issues such as spectral artifacts and reliance on carefully engineered substrates, limiting its versatility [11]. With its broad spectral range, AFM–IR faces challenges with low signal–to–noise ratio (SNR) and slow imaging speed [29]. O–PTIR also encounters limitations due to its narrow spectral range and slow imaging speed, further compounded by the very high acquisition cost [29].
@@
Nanoelectromechanical IR spectroscopy (NEMS–IR) is a technique based on the photothermal effect that has been introduced in recent years. It employs NEMS membrane resonators, or chips, as its key component [30, 31]. The NEMS chip can be used simultaneously as a sample carrier and detector. The NEMS chips, typically consisting of a pre–stressed silicon nitride (SiN) membrane, are robust and allow for a wide variety of sampling techniques [32], including direct aerosol collection [33, 31, 30] and drop casting. When a sample deposited on the NEMS chip’s surface absorbs IR light, local heating occurs, causing thermal expansion and tensile stress reduction. The corresponding frequency detuning of the NEMS chip is proportional to the absorbed power, and the frequency shift can be monitored using a closed–loop oscillation scheme in combination with a frequency counter, as demonstrated by Bešić et al. [34, 35]. NEMS–IR has shown picogram sensitivity at room temperature over a large spectral range, from the ultraviolet to the far–IR [32]. This technique has been used to analyze polymer nanoparticles [36, 37], pharmaceutical compounds [31, 38], polymer micelles [39], thin polymer films [40] and explosives [41]. NEMS–IR with in–situ thermal desorption (TD) analysis has also been demonstrated for separating simple analyte mixtures by Luhmann et al. [30].
NEMS–IR has only been used for dispersive spectroscopy, typically relying on expensive, narrowband tunable quantum cascade lasers (QCLs). This paper presents NEMS–FTIR spectroscopy as a new approach to the chemical characterization and quantification of nanoplastics. While cantilevers have previously been used as photothermal detectors in conjunction with FTIR sources [42], the approach presented here centers on NEMS membrane resonators shown in Fig. 1A, offering enhanced sensitivity and sampling flexibility. NEMS–FTIR uses a commercially available FTIR spectrometer as a light source, as shown in Fig. 1B, instead of a QCL. NEMS–FTIR, therefore, offers a powerful new approach for the routine monitoring and characterization of nanoplastics by combining the exceptional sensitivity and versatility of nanomechanical photothermal sensing with the broad spectral coverage and user–friendly operation of commercially available FTIR spectrometers. Here we demonstrate the power and versatility of this technique for the sampling, detection, characterization, and quantification of nanoplastics in aqueous samples and nylon tea bag leachates.
RESULTS
@@
Aqueous dispersions of three different polymers, PS (nominal diameter: 100 nm), PP (nominal diameter: 54 nm), and PVC (nominal diameter: 262 nm), were deposited via drop casting on individual NEMS chips. Fig. 2A–C shows PS, PP, and PVC nanoparticles after deposition onto the central circular perforation of the NEMS chips (see Fig. 1A). These scanning electron microscopy (SEM) images highlight the generally spherical morphology and relatively uniform size distributions of the nanoplastic particles studied. However, there are areas of nanoparticle agglomerations on the membrane of the NEMS chip caused by the coffee ring effect during droplet drying (see Fig. S1 in Supplementary Materials). After the nanodrop dispersion has dried, all nanoparticles are confined within the perforated area of the NEMS chip.
@@
Spectral data was obtained by depositing a 20 nL nanodrop of various nanoplastic concentrations on the NEMS chips (N=3 for each concentration). The NEMS–FTIR spectra of PS nanoplastic particles for each concentration, power–corrected, normalized using the SiN peak at 835 cm-1, and converted into absorbance, are shown in Fig. 2D. The details of the spectral processing procedure are described in the "Processing of NEMS–FTIR spectra" in the Materials and Methods section. The spectra of PP and PVC nanoparticles are available in the Supplementary Materials in Figs. S2 and S3, respectively.
@@
The IR peaks in the NEMS–FTIR spectra of PS nanoparticles correspond to data reported in the literature [43, 44, 45, 46]. Notably, PS is characterized by a mono–substituted benzene ring, as evidenced by a series of weak overtone and combination bands between 2000 and 1650 $\text{cm}^{-1}$ (so–called ’benzene fingers’), often difficult to detect in the spectrum if the concentration of molecules containing the benzene ring is low [45], and the peaks at 753 and 700 $\text{cm}^{-1}$. However, these peaks are clearly visible in the NEMS–FTIR spectra of PS, even for the very low masses in the nanogram range. The NEMS–FTIR spectra of PP and PVC nanoparticles also display peaks corresponding to their respective vibrational modes. The positions of these peaks are in accordance with prior studies using FTIR spectroscopy [45, 47], as well as with the measurements obtained using early IR spectrophotometers employing prisms [48]. NEMS–FTIR does not suffer from the red–shift typically associated with measurements performed by ATR–FTIR, as observed in previous studies [49]. In ATR–FTIR, slight shifts in some peak positions are observed due to differences in how this technique probes the sample, particularly the interaction of the evanescent wave with the sample surface and its dependence on refractive index and wavelength [50, 51].
NEMS–FTIR generates spectra comparable to those obtained with transmission–FTIR spectroscopy, allowing seamless integration with existing spectral libraries and chemometric tools for efficient and accurate analysis. Furthermore, since the nanoplastic particles are analyzed under high vacuum, the resulting spectra are free from spectral interferences from water molecules and carbon dioxide.
@@
Fig. 2E shows the calibration curves for PS, PP, and PVC nanoparticles. Calibration curves were constructed using the characteristic peaks at 1452 $\text{cm}^{-1}$ for PS (see inset in Fig. 2D), 1377 $\text{cm}^{-1}$ for PP (see inset in Fig. S2), and 1427 $\text{cm}^{-1}$ for PVC (see inset in Fig. S3) and the results showed good linearity ($\text{R}^{2}>0.9885$). LoDs were calculated from the calibration curves using the relation $3\sigma/m$, where $\sigma$ is the standard deviation of method blanks ($\text{N}=9$) and $m$ is the slope of the corresponding calibration curve. The results are summarized in Table 1. PP exhibited the steepest calibration slope, resulting from the strongest IR band used for the calibration curve, and consequently had the lowest LoD. Detection limits are lower than those typically achieved by Py–GC/MS [19] and are comparable to the performance of state–of–the–art TD–PTR–MS techniques [21]. The variations in detection limits for different nanoplastic particles did not correlate with nanoparticle size. At higher sample loadings, the observed nanoparticle agglomerations did not result in deviations in linearity.
@@
The mass of PS nanoparticles present on the NEMS chips was estimated using the attenuation coefficient of PS and the absorptances calculated from the NEMS–FTIR spectra at $\tilde{\nu}=1452\text{ cm}^{-1}$. The mass estimation procedure is detailed in the "Mass estimation" in the Materials and Methods section. The calculation of the sample absorptance accounted for the spatial distribution of the particles, which predominantly formed a circular patch in the center of the perforation on the NEMS chips. Fig. 3A shows the correlation between the deposited (amount of deposited PS particles) and estimated (from the measured absorptances of PS particles) mass of the nanoplastic particles. The uncertainty on the sample mass deposited on the NEMS chip was estimated from the uncertainty on the tools used for sample preparation and for sample deposition. The largest uncertainty on the sample mass as estimated from the NEMS–FTIR spectra arises from the variability in the measured absorptance $\alpha_{\text{S}}(1452\leavevmode\nobreak\ \text{cm}^{-1})$. Exact values for the masses deposited on the NEMS chips and estimated masses as calculated from the NEMS–FTIR spectra, along with their standard deviations, can be found in Table S1 in the Supplementary Materials.
@@
Fig. 3B presents a comparison between the NEMS–FTIR spectrum of 20 ng of PS, expressed in absorbance, and a reference spectrum calculated from refractive index data [52] for an equal sample mass and spot size showing agreement for both peak positions and intensity.
@@
Table 1: LoD calculation parameters. Summary of the calculated LoDs for different nanoplastics, including the corresponding wavenumbers, standard deviations of the blank ($\sigma$) at those wavenumbers (N = 9), and calibration slopes ($m$). LoDs were determined for $3\sigma/m$.
@@
Plastic Type Wavenumber (cm-1) Standard deviation of the blank ( $10^{-4}$ AU) Slope ($10^{-4}$ AU/ng) LoD (pg)
@@
The SEM images in Fig. 4A, B provide the close–up views of the perforated area of the NEMS chip sampled with PS, PP, and PVC nanoparticles previously mixed in a 1:1:1 mass ratio.
The NEMS–FTIR spectrum of the mixture in Fig. 4C shows the characteristic peaks of each polymer component, highlighted by vertical lines. Despite a mass of only 5 ng per component, the peaks corresponding to PS, PP, and PVC are clearly distinguishable. The exact peak positions of the PS, PP, and PVC nanoparticles can be found in the Supplementary Materials, section A. In environmental contexts, where mixtures of different polymer micro– and nanoparticles are commonly encountered, the broad spectral range of NEMS–FTIR increases the likelihood of detecting distinct peaks without interference, facilitating the identification of individual components when their spectral features are sufficiently resolved.
Figure 4: Characterization of nanoplastic mixture (A) SEM image of a mixture containing PS, PP, and PVC. (B) Magnified view of the membrane area highlighting PS (⌀ 100 nm, blue square), PP (⌀ 54 nm, green square), and PVC nanoparticles (⌀ 262 nm, red square). (C) Stacked spectra of pure PS, PP, and PVC nanoparticles and their 1:1:1 mixture. Vertical lines mark the characteristic peaks of each polymer.
@@
Fig. 5A shows the schematic of the experimental setup for the quantification of nylon tea bag leachates (TBL). The empty tea bag was immersed in 200 mL of preheated water at 95°$\degree$°C for 10 minutes.
@@
Figure 5: Characterization of PA particles released from nylon tea bags. (A) Schematic illustrating the sample preparation process. (B) SEM image showing fragments released into water by a nylon tea bag. (C) NEMS–FTIR spectra of nylon tea bag leachates (TBL, 100 nL aliquots deposited onto NEMS chips without pre–concentration) compared to a reference spectrum of the nylon tea bag (TB) recorded by ATR–FTIR (PerkinElmer, 4 cm-1 resolution, 16 scans). Vertical lines indicate characteristic IR peaks associated with PA. (D) NEMS–FTIR spectra of the leachates from a nylon tea bags subjected to accelerated weathering for 25 consecutive days, simulating one year of aging under adjusted conditions. The right inset graph highlights the characteristic PA peak at 1640 cm-1, corresponding to the amide I band. The left inset graph represents the dependence of signal intensity at 1640 cm-1 on the number of days during which the nylon tea bag was subjected to accelerated aging.

### v2

# Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR

@@ Abstract
We present a photothermal infrared spectroscopy-based approach for the chemical characterization and quantification of nanoplastics. By combining the high sensitivity of nanoelectromechanical systems (NEMS) with the wide spectral range and ubiquity of commercially available Fourier transform infrared (FTIR) spectrometers, NEMS–FTIR offers a cost-efficient and cryogen-free option for the rapid, routine analysis of nanoplastics in aqueous samples. Polypropylene, polystyrene, and polyvinyl chloride nanoplastics with nominal diameters ranging from 54 to 262 nm were analyzed by NEMS–FTIR with limits of detection ranging from 102 pg to 355 pg. The absorptance measured by NEMS–FTIR could be further converted to absolute sample mass using the attenuation coefficient, as demonstrated for polystyrene. Thanks to the wide spectral range of NEMS-–FTIR, nanoplastic particles from different polymers could be readily identified, even when present in a mixture. The potential of NEMS–FTIR for the analysis of real samples was demonstrated by identifying the presence of nanoplastics released in water during tea brewing. Polyamide leachates were identified in the brewing water without sample pre–concentration. Accelerated aging of the tea bags under elevated temperature and UV radiation showed further release of polyamide over time.
@@ INTRODUCTION
Nanoplastics have become ubiquitous and pose significant environmental and health risks due to their high reactivity, potential to transport pollutants, and widespread dispersion across ecosystems [1, 2, 3, 4, 5, 6, 7]. The ability of nanoplastics to penetrate deep into tissues underscores the need for their routine chemical characterization and monitoring [8]. Sampling of nanoplastic particles is particularly challenging due to their nanoscale dimensions and low concentrations, which often exceed the capabilities of many analytical techniques [9, 10, 11]. While state–of–the–art techniques allow for routine microplastic detection and identification, nanoplastics, on the other hand, remain a challenge to this day.
@@
Mass spectrometry–based techniques can be used for the bulk chemical identification of nanoplastics [10, 11, 12]. Pyrolysis–gas chromatography–mass spectrometry (Py–GC/MS) has been extensively applied to detect various types of nanoplastics, including polystyrene (PS), polypropylene (PP), polycarbonate (PC), polyethylene terephthalate (PET), polymethyl methacrylate (PMMA), polyvinyl chloride (PVC), polyethylene (PE), and Nylon 6/66, with reported limits of detection (LoDs) as low as 1 ng/L [13, 14]. Using the same technique, Li et al[15] estimated an LoD of 10 ng for two polymer types of nanoplastics in water, PS, and PMMA. A combination of atomic force microscope–IR (AFM–IR) spectroscopy and Py–GC/MS was also used to detect and quantify PE and PVC particles (20-–1000 nm), with LoDs of 0.28 $\mu$g for PE and 1.54 $\mu$g for PVC [16]. Alternative approaches, such as thermal desorption–proton transfer reaction-mass spectrometry (TD–PTR–MS), demonstrated remarkable sensitivity, with Materić et al[17] detecting pure 1 $\mu$m PS particles with an LoD of 0.34 ng, allowing detection of these particles in environmental samples as small as 1 mL without preconcentration. While mass spectrometry–based techniques provide excellent capabilities, their complexity and cost hinder their use in routine environmental monitoring [11, 12, 18].
FTIR spectroscopy with its broad spectral range allows functional group chemical identification of unknown samples and is one of the most extensively used methods in microplastic analysis [19, 20]. Commonly used attenuated total reflectance–FTIR (ATR–FTIR) spectroscopy is a contact bulk analysis technique with a detection limit in the order of 10 ng [21]. FTIR microscopy methods, such as µ–FTIR scanning microscopy and focal plane array–FTIR (FPA–FTIR) microscopy, offer a contact–less FTIR solution, reaching LoDs as low as 450 pg when used with a cryogenically cooled MCT array [22]. However, diffraction limits the reliable analysis of single particles with sizes below 10 $\mu$m [19, 20].
Quantum cascade laser infrared (QCL–IR) microscopy provides higher sensitivity, albeit with a narrower spectral range and additional issues with coherence artifacts. Although QCLs cover the fingerprint region relevant for polymer identification and provide stronger spectral signals due to their high source intensity, they typically have higher noise levels and limited spectral coverage, missing informative regions such as the C–-H stretching vibrations around 3000 cm-1 and the low-wavenumber region around 700 cm-1. Full–range spectra, as enabled by FTIR, also allow for chemometric analysis and spectral deconvolution, which can help identify and quantify nanoplastics in multicomponent samples. While such methods can also work within narrower spectral windows, having access to broader spectral information can make differentiation easier and more reliable. QCL–IR microscopy is a complementary method to FTIR microscopy and is therefore sometimes offered in tandem. [^2] HYPERION II from Bruker Corporation All standard IR spectroscopy techniques suffer from Mie scattering–related spectral artifacts when measuring particulate samples.[23]
Scattering, diffraction, and coherence issues can be mitigated by QCL–based photothermal scanning methods, such as optical–-photothermal IR (O-PTIR) spectroscopy or AFM-IR microscopy. O–PTIR has proven effective for the imaging with sub–micrometer resolution of individual PS nanoparticles as small as 250 nm in mammalian tissues [24] and identification of plastics (> 600 nm) containing polysiloxanes and imides, originating from the steam disinfection of silicone baby teats [25]. Even better lateral resolution and single nanoparticle analysis can be obtained by AFM-IR [26]. However, AFM–IR faces challenges with low signal–to–noise ratio (SNR) and slow imaging speed [19].
Raman spectroscopy has the advantage of probing at visible wavelengths with low interference from water and, similar to O–PTIR, allows for the analysis of plastic particles down to sub-micrometer sizes, with spatial resolution sufficient to resolve and chemically identify individual particles. [27]. Surface–enhanced Raman spectroscopy (SERS) has emerged as a promising technique. While Zhou et al[28] and Hu et al[29] reported LoDs as low as 5 $\mu$g/mL for PS particles as small as 50 nm using conventional SERS substrates, integration with sample preparation methods has pushed the sensitivity down to $10^{-4}$ $\mu$g/mL [30]. Hyperspectral stimulated Raman scattering (SRS) microscopy has been demonstrated as a powerful tool for the chemical imaging of nanoplastics, enabling identification and classification of particles down to 130 nm [31]. However, SERS suffers from issues such as spectral artifacts and reliance on carefully engineered substrates, limiting its versatility [11, 32]. SRS microscopy, despite offering rapid detection and high sensitivity, relies on complex and expensive instrumentation [33].
@@
Although existing analytical techniques such as Py–GC/MS, QCL–IR, O–PTIR, or SERS offer high sensitivity and specificity, they remain impractical for routine environmental monitoring, particularly for water suppliers and official control laboratories [34].
@@
Nanoelectromechanical systems IR spectroscopy (NEMS–IR) is a technique based on the photothermal effect that has been introduced in recent years. It employs NEMS membrane resonators, or chips, as its key component [35, 36]. The NEMS chip can be used simultaneously as a sample carrier and detector. The NEMS chips, typically consisting of a pre–stressed few–nanometer–thin silicon nitride (SiN) membrane, are robust and allow for a wide variety of sampling techniques [37], including direct aerosol collection [38, 36, 35] and drop casting. When a sample deposited on the NEMS chip’s surface absorbs IR light, local heating occurs, causing thermal expansion and tensile stress reduction. The corresponding frequency detuning of the NEMS chip is proportional to the absorbed power, and the frequency shift can be monitored using a closed–loop oscillation scheme in combination with a frequency counter, as demonstrated by Bešić et al[39, 40]. NEMS–IR has shown picogram sensitivity at room temperature over a large spectral range, from the ultraviolet to the far–IR [37]. This technique has been used to analyze polymer nanoparticles [41, 42], pharmaceutical compounds [36, 43], polymer micelles [44], thin polymer films [21] and explosives [45]. NEMS–IR with in–situ thermal desorption (TD) analysis has also been demonstrated for separating simple analyte mixtures by Luhmann et al[35].
NEMS–IR has only been used for dispersive spectroscopy, typically relying on narrow-band tunable QCLs. This paper introduces NEMS–FTIR spectroscopy (see Figure 1) as a new approach to the chemical characterization and quantification of nanoparticles, demonstrated here with nanoplastics. By combining the broad spectral coverage, wide availability, and ease of use of commercially available FTIR spectrometers with the high sensitivity of nanomechanical photothermal detection using NEMS membrane resonators, NEMS-–FTIR enables artifact–free measurements with picogram–level detection limits — all without the need for cryogenic cooling. Thanks to its straightforward sampling of nanoparticles dispersions directly onto the disposable NEMS chip, NEMS-–FTIR provides a powerful bulk analysis method for the routine detection and characterization of nanoplastics. Optomechanical metalized cantilevers have previously been used as photothermal detectors in combination with FTIR sources [46], but their lower sensitivity [47], reliance on optical readout, and limited surface area for sample collection have constrained their practical application.
Here we demonstrate the capability of NEMS–FTIR for the sampling, identification, and quantification of nanoplastics dispersions and nylon-based polyamide (PA) tea bag leachates. Many studies have analyzed the release of micro- and nanoplastic particles from tea bags during tea preparation using different analytical methods such as scanning transmission X–ray microscopy [48], IR photothermal heterodyne microscopy [49], Raman imaging [50], and ATR–FTIR spectroscopy [51, 52]. Most of these, however, require a much higher tea bag–to–water ratio than that applied in this study [52, 51] or involve sample up–concentration steps such as ultracentrifugation [51], density separation [48], or evaporation drying to obtain a powder [52]. In contrast, we demonstrate here that NEMS-–FTIR can chemically identify PA leachates in the brewing water from a single tea bag. NEMS-FTIR analysis of the leachate required just a few nL sample aliquot and did not require any pre–concentration steps.
RESULTS AND DISCUSSION
@@
Aqueous dispersions of three different polymers, PS (nominal diameter: 100 nm), PP (nominal diameter: 54 nm), and PVC (nominal diameter: 262 nm), were deposited via drop casting on individual NEMS chips. Figure 2A–C shows PS, PP, and PVC nanoparticles after deposition onto the central circular perforation of the NEMS chips (see Figure 1A). These scanning electron microscopy (SEM) images highlight the generally spherical morphology and relatively uniform size distributions of the nanoplastic particles studied. However, there are areas of nanoparticle agglomerations on the membrane of the NEMS chip caused by the coffee ring effect during droplet drying (see Figure S1 in Supporting Information). After drying, nanoparticles were confined within the perforated area of all analyzed NEMS chips, except in a few instances where minor droplet spreading occurred beyond this region. The spread–out particles still remained located within the confines of the NEMS membrane, which is the active area relevant for signal detection.
@@
Spectral data were obtained by depositing a 20 nL drop of various nanoplastic concentrations on the NEMS chips (N=3 for each concentration). The NEMS–FTIR spectra of PS nanoplastic particles for each concentration, power–corrected, normalized using the SiN peak at 835 cm-1, and converted into absorbance, are shown in Figure 2D. The broad SiN peak, originating from the membrane material and spanning the 1000–700 cm-1 region, serves as a useful internal standard for spectral normalization. The details of the spectral processing procedure are described in "Processing of NEMS–FTIR spectra" in the Materials and Methods section. The spectra of PP and PVC nanoparticles are available in the Supporting Information in Figures S2 and S3, respectively.
@@
The IR peaks in the NEMS–FTIR spectra of PS nanoparticles are in agreement with data reported in the literature [53, 54, 55, 56]. Notably, PS is characterized by peaks at 753 and 700 $\text{cm}^{-1}$, as well as by a mono–substituted benzene ring, as evidenced by a series of weak overtone and combination bands between 2000 and 1650 $\text{cm}^{-1}$ (so–called ’benzene fingers’), often difficult to detect in the spectrum if the concentration of molecules containing the benzene ring is low [55]. However, these peaks are clearly visible in the NEMS–FTIR spectra of PS, even for the very low masses in the nanogram range.
The NEMS–FTIR spectra of PP and PVC nanoparticles also display peaks which are in agreement with prior studies using FTIR spectroscopy [55, 57], as well as with the measurements obtained using early IR spectrophotometers employing prisms [58]. NEMS–FTIR does not suffer from the red shift typically associated with measurements performed by ATR–FTIR, as observed in previous studies [59]. In ATR–FTIR, slight shifts in some peak positions are observed due to the interaction of the evanescent wave with the sample surface, which depends on both the refractive index and wavelength. The wavelength-dependent penetration depth inherent to ATR measurements can also cause lower–wavenumber bands to appear more intense compared to transmission (or absorbance) spectra. [60, 61, 62, 63, 64]. This artefact is also absent from NEMS–FTIR spectra. ATR–FTIR performance is inherently constrained by the choice of crystal material, as different sample types require different ATR crystals for optimal spectral response [63]. This issue also does not apply to NEMS-FTIR, which relies on direct absorption. NEMS–FTIR generates spectra comparable to those obtained with transmission–FTIR spectroscopy, allowing seamless integration with existing spectral libraries and chemometric tools for efficient and accurate analysis. Furthermore, since the nanoplastic particles are analyzed under high vacuum, the resulting spectra are free from spectral interferences from water molecules and carbon dioxide (CO2).
@@
Figure 2E shows the calibration curves for PS, PP, and PVC nanoparticles. Calibration curves were constructed using the characteristic peaks at 1452 $\text{cm}^{-1}$ for PS (see inset in Figure 2D), 1377 $\text{cm}^{-1}$ for PP (see inset in Figure S2), and 1427 $\text{cm}^{-1}$ for PVC (see inset in Figure S3) and the results showed good linearity ($\text{R}^{2}>0.9885$). LoDs were calculated from the calibration curves using the relation $3\sigma/m$, where $\sigma$ is the standard deviation of method blanks ($\text{N}=9$) and $m$ is the slope of the corresponding calibration curve. The results are summarized in Table 1. PP exhibited the steepest calibration slope, resulting from the strongest IR band used for the calibration curve, and consequently had the lowest LoD. Detection limits are lower than those typically achieved by Py–GC/MS [15] and are comparable to the performance of state–of–the–art TD–PTR–MS techniques [17]. The variations in detection limits for different nanoplastic particles did not correlate with nanoparticle size. In the measurements conducted on the chips with higher loadings, the observed nanoparticle agglomerations did not result in deviations in linearity.
@@
The mass of PS nanoparticles present on the NEMS chips was estimated using the attenuation coefficient of PS and the absorptances calculated from the NEMS–FTIR spectra at $\tilde{\nu}=1452\text{ cm}^{-1}$. The mass estimation procedure is detailed in the "Mass estimation" in the Methods section. The calculation of the sample absorptance accounted for the spatial distribution of the particles, which predominantly formed a circular patch in the center of the perforation on the NEMS chips. Figure 3A shows the correlation between the deposited (amount of deposited PS particles) and estimated (from the measured absorptances of PS particles) mass of nanoplastic particles. The uncertainty on the sample mass deposited on the NEMS chip was estimated from the uncertainty on the tools used for sample preparation and for sample deposition. The largest uncertainty on the sample mass as estimated from the NEMS–FTIR spectra arises from the variability in the measured absorptance $\alpha_{\text{S}}(1452\leavevmode\nobreak\ \text{cm}^{-1})$. Exact values for the masses deposited on the NEMS chips and estimated masses as calculated from the NEMS–FTIR spectra, along with their standard deviations, can be found in Table S1 in the Supporting Information.
@@
Figure 3B presents a comparison between the NEMS-FTIR spectrum of 20 ng of PS, expressed in absorbance, and a reference spectrum calculated from refractive index data [65] for an equal sample mass and spot size, showing agreement for both peak positions and intensity.
@@
Table 1: LoD calculation parameters. Summary of the calculated LoDs for different nanoplastics, including the corresponding peak position (in wavenumbers), standard deviations of the blank ($\sigma$) at those respective peak positions (N = 9), and calibration slopes ($m$). LoDs were determined for $3\sigma/m$.
@@
Plastic Type Peak position (cm-1) Standard deviation of the blank ( $10^{-4}$ AU) Slope ($10^{-4}$ AU/ng) LoD (pg)
@@
While NEMS–FTIR can be used to detect very small masses of analytes, very dilute solutions do require pre–concentration. The largest droplet volume that can be drop cast on a NEMS chip is limited to approximately 500 nL. Considering an LOD of 353 pg for PS (see Table 1), the minimum detectable concentration of PS is 0.7 $\mu$g/mL for a single 500 nL drop of sample. However, in addition to usual pre–concentration techniques such as solid phase extraction, ultrafiltration, and diafiltration, the NEMS–FTIR chip allows new sample pre–concentration possibilities. Multiple droplets can be deposited consecutively on the NEMS chip with intermediate drying steps to increase the analyte mass deposited to the desired level. Sampling via nebulization can also be used to concentrate larger sample volumes onto the NEMS chip. Based on a typical nebulizer flow rate of 20 $\mu$L/min, a sampling time of 1 hour, and an aerosol capture efficiency of approximately 50 % for 100 nm particles [38], we estimate that concentrations as low as 0.6 ng/mL could be measurable. When dealing with complex matrices containing high concentrations of organic material relative to the target analyte, matrix simplification or removal may be required to remove spectral interferences or avoid overloading the NEMS chip.
@@
The SEM images in Figure 4A, B provide the close–up views of the perforated area of the NEMS chip sampled with PS, PP, and PVC nanoparticles previously mixed in a 1:1:1 mass ratio.
The NEMS–FTIR spectrum of the mixture in Figure 4C shows the characteristic peaks of each polymer component, highlighted by vertical lines. Despite a deposited mass of only 5 ng per component, the peaks corresponding to PS, PP, and PVC are clearly distinguishable. The exact peak positions of the PS, PP, and PVC nanoparticles can be found in the Supporting Information, section A. In environmental contexts, where mixtures of different polymer micro- and nanoparticles are commonly encountered, the broad spectral range of NEMS–FTIR increases the likelihood of detecting distinct peaks without interference, facilitating the identification of individual components when their spectral features are sufficiently resolved.
Figure 4: Characterization of nanoplastic mixture. (A) SEM image of a mixture containing PS, PP, and PVC. (B) Magnified view of the membrane area highlighting PS (⌀ 100 nm, blue square), PP (⌀ 54 nm, green square), and PVC nanoparticles (⌀ 262 nm, red square). (C) Stacked NEMS–FTIR spectra of pure PS, PP, and PVC nanoparticles and their 1:1:1 mixture. Vertical lines mark the characteristic peaks of each polymer.
@@
Figure 5A shows the schematic of the experimental setup for the quantification of nylon tea bag leachates. The empty tea bag was immersed in 200 mL of water preheated at 95°$\degree$°C for 10 minutes.
@@
Figure 5: Characterization of PA particles released from nylon tea bags. (A) Schematic illustration of the sample preparation process. (B) SEM image showing fragments released into water by a nylon tea bag. (C) NEMS–FTIR spectra of nylon tea bag leachates (TBL; 100 nL, and 500 nL aliquots deposited onto NEMS chips without pre–concentration) compared to ATR–FTIR spectrum of 500 nL of tea bag leachate (TBL) and the ATR–FTIR spectrum of the nylon tea bag (TB). Vertical lines indicate characteristic IR peaks associated with PA. (D) NEMS–FTIR spectra of the leachates from nylon tea bags subjected to accelerated weathering for 25 consecutive days, simulating one year of aging under adjusted conditions. The right inset graph highlights the characteristic PA peak at 1640 cm-1, corresponding to the amide I band. The left inset graph represents the dependence of signal intensity at 1640 cm-1 on the number of days during which the nylon tea bag was subjected to accelerated aging.